# Language Models Surface the Unwritten Code of Science and Society

### v4

@@ Abstract
This paper calls on the research community not only to investigate how human biases are inherited by large language models (LLMs) but also to explore how these biases in LLMs can be leveraged to make society’s “unwritten code” — such as implicit stereotypes and heuristics — visible and accessible for critique. We introduce a conceptual framework through a case study in science: uncovering hidden rules in peer review - the factors that reviewers care about but rarely state explicitly due to normative scientific expectations. The idea of the framework is to push LLMs to speak out their heuristics through generating self-consistent hypotheses - why one paper appeared stronger in reviewer scoring - among paired papers submitted to 45 academic conferences, while iteratively searching deeper hypotheses from remaining pairs where existing hypotheses cannot explain. We observed that LLMs’ normative priors about the internal characteristics of good science extracted from their self-talk, e.g., theoretical rigor, were systematically updated toward posteriors that emphasize storytelling about external connections, such as how the work is positioned and connected within and across literatures. Human reviewers tend to explicitly reward aspects that moderately align with LLMs’ normative priors (correlation = 0.49) but avoid articulating contextualization and storytelling posteriors in their review comments (correlation = –0.14), despite giving implicit reward to them with positive scores. These patterns are robust across different models and out-of-sample judgments. We discuss the broad applicability of our proposed framework, leveraging LLMs as diagnostic tools to amplify and surface the tacit codes underlying human society, enabling public discussion of revealed values and more precisely targeted responsible AI.
Language Models Surface the Unwritten Code of Science and Society
@@ 1 Introduction
Large language models (LLMs), extensively trained on vast corpora of human-generated text and visuals, inevitably absorb and reflect the patterns of human thought embedded within these datasets. Researchers have explored how AI can be used to judge human activities such as scientific research [37], misinformation oversight [12], and perceptions of physical appearance [14]. With strong predictive capacities, LLMs can replicate shifts in public opinion and human beliefs on contentious political issues, including same-sex marriage [26] and climate change [34]. These models often reveal biases that closely align with directly measured human biases and even amplify them across various social dimensions [73], such as gender [27], race [23], social identity [24], and profession [21]. Such biases are not computational artifacts but reflect the deeper cultural and social frameworks that shape human communication [6, 33, 19]. The recognition that LLMs encapsulate and propagate societal biases and heuristics is now well-established in literature, framing these models as inherently cultural and social technologies [11, 16, 65, 46, 35]. This paper moves beyond merely acknowledging LLM biases; it proposes leveraging them as diagnostic tools to uncover and critically examine the implicit, unwritten societal codes that govern human interaction.
@@
The “unwritten code" – comprising tacit norms, implicit stereotypes, and subtle conventions that invisibly influence societal dynamics – represents a potent force that, if left unexamined, unsummarized, and unaddressed, can perpetuate systemic inequities and injustices. Leveraging the inherent biases embedded within LLMs can shed light on otherwise elusive patterns that sustain societal inequality. One straightforward path to mitigating human bias involves summarizing and clearly articulating the biases humans hold, subsequently enabling evaluation, regulation and systemic change. Nevertheless, these biases frequently exist "unwritten" and thus resist formalization and systematic analysis. Even though some methods can expose bias in LLMs when models appear to avoid expressing it, as by analyzing associated words [7], there remains a deeper layer of unwritten code that contemporary simple techniques cannot recover. These deeper structures activate during the model’s reasoning process, serving as heuristics that streamline its efficiency  [^1] We have more discussions in Section 6 and Appendix G.. An algorithmic framework capable of extracting and explicitly articulating them through self-reinforcement and amplification of the rationale within LLMs, which are inherently imbued with human biases and heuristics, could illuminate the content and character of bias within models and human cultures.
Thus, this paper calls for an intentional shift in perspective: We call on researchers not only to investigate how human biases are inherited by LLMs but also to explore how these biases can be leveraged to make the unwritten codes visible, as shown in Figure 1. To this end, we present a conceptual framework that enables this process by pushing LLMs to speak out their heuristics through searching their deeper self-consistent hypotheses that can explain their decision-making. By prompting models to think out of the box and "exaggerate" hidden rules that humans might be unable or unwilling to acknowledge openly, we can more clearly illustrate how such biases and heuristics operate and which groups they potentially disadvantage.
@@ 2 A Case Study: The Aesthetic of Peer Review
What makes a scientific paper "good"? While scientific communities often emphasize norms such as rigor, the actual process of evaluating quality is far more entangled with social dynamics, implicit biases, and tacit expectations [8]. Peer review, the central gatekeeping mechanism of science, is deeply shaped by these social dynamics. Despite its appearance of objectivity, peer review can be biased toward well-known institutions, established authors, or fashionable topics [31]. Reviewers may assign high scores to work that aligns with unspoken disciplinary tastes or collective values [57], while "lazily" masking these preferences in the language of standardized evaluation criteria [48]. As a result, much of what constitutes the tacit code of a field – its evaluative culture – remains hidden beneath the surface, shaping scientific recognition in nontransparent ways that remain hard to alter.
@@
Data: We collected data from the OpenReview API and Paper Copilot  [^2] https://papercopilot.com/, a website that aggregates peer review scores (mostly) for computer science conference submissions. Our final dataset includes metadata, peer review scores, and review comments for 26,731 papers submitted to 45 conferences including computer science, physics, medicine, and social sciences. We use peer review scores as a proxy for the paper’s perceived quality, as shown in Appendix A, due to the consistency and high confidence of judgments among reviewers.
@@
To represent each paper’s content, we follow the approach of [71], constructing an extended abstract that integrates additional structured elements — contextual background, key ideas, methodological and theoretical details, experiments and results, and mentioned impact. We generate these extended abstracts using prompts used in [71] with GPT-4o. This representation generation has been empirically shown to achieve high quality across evaluation metrics and to closely match human annotations. It also demonstrates strong alignment with the original text and a low rate of hallucination [4, 71], see Appendix E for fidelity experiments. We deliberately avoid feeding full texts directly to LLMs for two main reasons. First, context windows remain limited, preventing broad coverage that would enable more generalizable hypotheses. Current approaches, e.g., [45], employ reduction methods such as sparse autoencoders to compress long inputs into activated neurons for hypothesis generation, but these are typically applied only to straightforward cases (e.g., political attitudes) where activations can be reduced to obvious dimensions (e.g., gender and race). Second, prior work has shown that LLMs often struggle with long, unstructured inputs, leading to degraded performance and increased hallucination rates mid-text [41, 68]. When presented with lengthy texts, attention may diffuse across irrelevant sections, impairing reasoning over core content. To address these issues, we adopt a curated and condensed input format, ensuring that the model attends to the most relevant information. Dataset and paper representations are detailed in Appendix H. Most importantly, robustness checks in the Appendix E show that LLM judgments on full papers and extended abstracts yield highly consistent results (consistency = 0.9).
@@ 3 The Framework: Prior, Posterior, and Hypothesis Generation
We adopt a hypothesis generation framework – a topic increasingly explored in the automation of scientific discovery using AI [49, 40, 9, 39] – to update LLMs’ stereotyped priors about good science into posteriors that reflect how human evaluators and LLMs actually judge.
@@
Self-consistency Refinement: We apply the five generated hypotheses to all pairs in the dataset. For each pair and hypothesis, we ask the LLM to determine the degree to which the hypothesis explains the observed quality difference (human review scores) using the following prompt: "Analyze the content from both papers to determine whether Paper 2 is superior in the aspects identified in the hypothesis (i.e., if the comparison of content supports the hypothesis)." This type of self-confirmation explicitly reveals the inherent biases of LLMs during the judgment process. The LLM outputs two things: (1) a binary label (0 or 1) indicating whether the hypothesis explains the score difference, and (2) a confidence score (0–10) for its judgment. We use 3-fold voting and compute the final label using a confidence-weighted scheme: if the sum of confidence scores for the label "1" exceeds that of label "0", the final label is 1; otherwise, it is 0. This confidence-weighted approach has been shown to improve self-consistency and accelerate convergence to the stable answer of LLMs across a variety of QA tasks [58]. Beyond improving reliability, confidence scores across multiple judgments provide a fine-grained signal that distinguishes borderline cases from clear-cut ones - an especially valuable property in our setting, where evaluating internal characteristics is inherently difficult [8] and confidence levels can help identify when the model is genuinely uncertain [58]. We also tested majority voting, and the resulting labels are highly correlated with the confidence-weighted labels (correlation > 0.98). LLMs may exhibit position bias, meaning that when serving as judges, they tend to favor the candidate appearing in a specific position (e.g., the first) [53]. We adopt standard strategies to mitigate this bias. First, we construct paper pairs with substantial quality differences; prior research has shown that position bias is the most pronounced when the gap is small [36, 53]. Second, we randomize the positions of Paper 1 and Paper 2 across voting while maintaining logically consistent prompts and voting aggregation. Consistency of judgments across randomized positions achieves 77% and we have more discussions in Appendix M.
@@
We are particularly interested in belief change during this process, i.e., how LLM’s prior about good science transforms into a robust posterior, after judging the actual pairwise dataset. For each generated hypothesis, when the LLM judges the pairwise dataset, we retain the proportion of cases it can explain (one pair can be explained by multiple hypotheses). We also match each hypothesis to priors (prompting: do you think they are talking about the same idea and same hypothesis?) to obtain their appearance frequency in the prior. We also conducted cross-validation experiments for the reliability of LLM labels of "match" using a heuristic dictionary-based method (locating some keywords of posterior hypotheses in priors to identify "match"), and these two methods exhibit a relatively high degree of correlation of 0.7, pp=0.001 (see Appendix K for details). The frequency of each posterior hypothesis in the prior distribution is measured by counting how often a similar hypothesis appears within each window of 20 sequential prior hypotheses across 250 simulations. We use 20-hypothesis windows to align with the posterior generation setup, where LLMs are prompted to produce 4 rounds of 5 unique hypotheses. This ensures a fair comparison between the frequency of hypotheses in the prior and posterior distributions.
@@ 4.2 Human judgments and LLMs share priors, but not posteriors
We observe notable shifts in the frequency of certain hypotheses from prior to posterior, as shown in Figure 4. Some hypotheses remain relatively stable in frequency and their explanatory power for good science (middle). In contrast, others manifest substantial change – some decrease markedly in frequency (left), while others increase significantly (right) from the prior distribution (see Table 1). This shift reflects a transition, after being exposed to the pairwise data, from emphasizing normative scientific values – such as theoretical or methodological rigor - to prioritizing more practical, communicative, and contextualization aspects, including storytelling quality and interdisciplinary relevance. When treated as a peer reviewer, the LLM initially aligns with conventional scientific ideals. In practice, however, it leans more heavily on accessible, narrative-driven criteria that appear more explanatory in distinguishing lower from higher rated papers (Figure 3 panels c and d). This suggests a divergence between explicit and implicit criteria of evaluation in science, a phenomenon widely appreciated although rarely measured in social studies of science [30].
@@
We collected human review comments and annotated them using in-context learning prompts with GPT-4o, leveraging its strong performance in social computing annotation tasks, often surpassing human experts and crowd-workers [61, 1, 22]. For each hypothesis, the annotation model was provided with three illustrative examples of reviewers not-mentioning (label 0) specific aspects in a hypothesis, along with the reviewers’ attitudes toward those aspects (positive: 1 or negative: -1, they are considered as "mentioning"). For robustness, we conducted two experiments: 1) the dictionary method produces high consistency with LLM’s "mention" label with correlation of 0.78, $p=0.000$, detailed in Appendix K. 2) Three master’s students specializing in social computing independently conducted the human annotations, achieving a high level of consistency with the LLM-generated labels (see Appendix B for details). We find human comments align with LLM’s priors in terms of the frequency of aspects they address, meaning that they articulate a similar attention distribution as LLM’s prior. In contrast, human reviewers rarely articulate their updated judgments like the LLM does. That is, if we assume the LLM mirrors human reviewing behavior, both begin from similar priors. However, while the LLM transparently updates its evaluation (posterior), human reviewers often refrain from articulating such shifts – likely due to normative expectations in academic reviewing – despite favoring unarticulated values with higher scores (we have more discussions and empirical evidence in Section 6).
@@
5 Generalizability
We briefly discuss examples to generalize our framework to other domains such as hiring, education, medicine, and criminal justice.
Resume Screening: In hiring, LLMs could be tasked with ranking pairs of resumes. Initially, their evaluations might focus on normative criteria such as educational background, work experience, and relevant skills. However, by pushing the models to elaborate beyond surface-level assessments, we can prompt them to elevate less normative but influential factors – such as perceived cultural fit – which account for differences in longer-term employee performance and retention. For example, an LLM might infer that a candidate who has previously worked at firms different in culture from the hiring company (e.g., Google vs. Amazon) is not a good fit, even if this is not explicitly stated in job descriptions [50].
College Admission: Similar to our main study, LLMs can be used to judge student statements/college admission essays while being encouraged to articulate implicit reasoning. Prior work suggests that essays written by students from certain ethnic or linguistic backgrounds often exhibit distinct writing styles, which can inadvertently be recognized and penalized by human graders [32, 2, 3]. By analyzing patterns in LLMs’ scoring rationales, we can investigate whether such signals contribute to systematic biases in college admission that undermine educational equity.

### v5

# Position: Language Models Should be Used to Surface the Unwritten Code of Science and Society

@@ Abstract
This position paper calls on the research community not only to investigate how human biases are inherited by large language models (LLMs) but also to explore how these biases in LLMs can be leveraged to make society’s "unwritten code" — such as implicit stereotypes and heuristics — visible and accessible for critique. We introduce a conceptual framework through a case study in science: uncovering hidden rules in peer review - the factors that reviewers care about but rarely state explicitly due to normative scientific expectations. The idea of the framework is to push LLMs to speak out their heuristics through generating self-consistent hypotheses - why one paper appeared stronger in reviewer scoring - among paired papers submitted to 46 academic conferences, while iteratively searching deeper hypotheses from remaining pairs where existing hypotheses cannot explain. We observed that LLMs’ normative priors about the internal characteristics of good science extracted from their self-talk, e.g., theoretical rigor, were systematically updated toward posteriors that emphasize storytelling about external connections, such as how the work is positioned and connected within and across literatures. Human reviewers tend to explicitly reward aspects that moderately align with LLMs’ normative priors (correlation = 0.49) but avoid articulating contextualization and storytelling posteriors in their review comments (correlation = –0.14), despite giving implicit reward to them with positive scores. These patterns are robust across different models and out-of-sample judgments. We discuss the broad applicability of our proposed framework, leveraging LLMs as diagnostic tools to amplify and surface the tacit codes underlying human society, enabling public discussion of revealed values and more precisely targeted responsible AI.
Position: Language Models Should be Used to Surface
the Unwritten Code of Science and Society
@@ 1 Introduction
Large language models (LLMs), extensively trained on vast corpora of human-generated text and visuals, inevitably absorb and reflect the patterns of human thought embedded within these datasets. Researchers have explored how AI can be used to judge human activities such as scientific research (Liang et al., 2024), misinformation oversight (Chen et al., 2024a), and perceptions of physical appearance (Conwell et al., 2025). With strong predictive capacities, LLMs can replicate shifts in public opinion and human beliefs on contentious political issues, including same-sex marriage (Kim and Lee, 2023) and climate change (Lee et al., 2024b). These models often reveal biases that closely align with directly measured human biases and even amplify them across various social dimensions (Ziems et al., 2024). While this position paper moves beyond merely acknowledging LLM biases; it calls on leveraging them as diagnostic tools to uncover and examine the implicit, unwritten societal codes that govern human interaction.
@@
The “unwritten code" – comprising tacit norms, implicit stereotypes, and subtle conventions that invisibly influence societal dynamics – represents a potent force that, if left unexamined, unsummarized, and unaddressed, can perpetuate systemic inequities and injustices. Leveraging the inherent biases embedded within LLMs can shed light on otherwise elusive patterns that sustain societal inequality. One straightforward path to mitigating human bias involves summarizing and clearly articulating the biases humans hold, subsequently enabling evaluation, regulation and systemic change. Nevertheless, these biases frequently exist "unwritten" and thus resist formalization and systematic analysis. Even though some methods can expose bias in LLMs when models appear to avoid expressing it, as by analyzing associated words (Bai et al., 2025), there remains a deeper layer of unwritten code that contemporary simple techniques cannot recover. These deeper structures activate during the model’s reasoning process, serving as heuristics that streamline its efficiency  [^1] We have more discussions in Section 5 and Appendix G.. An algorithmic framework capable of extracting and explicitly articulating them through self-reinforcement and amplification of the rationale within LLMs, which are inherently imbued with human biases and heuristics, could illuminate the content and character of bias within models and human cultures. Thus, this position paper calls for an intentional shift in perspective: We call on researchers not only to investigate how human biases are inherited by LLMs but also to explore how these biases can be leveraged to make the unwritten codes visible, as shown in Figure 1. To this end, we present a conceptual framework that enables this process by pushing LLMs to speak out their heuristics through searching their deeper self-consistent hypotheses that can explain their decision-making. By prompting models to think out of the box and "exaggerate" hidden rules that humans might be unable or unwilling to acknowledge openly, we can more clearly illustrate how such biases and heuristics operate and which groups they potentially disadvantage.
@@ 2 A Case Study: The Aesthetic of Peer Review
What makes a scientific paper "good"? While scientific communities often emphasize norms such as rigor, the actual process of evaluating quality is far more entangled with social dynamics, implicit biases, and tacit expectations (Bao and Teplitskiy, 2024). Peer review, the central gatekeeping mechanism of science, is deeply shaped by these social dynamics. Despite its appearance of objectivity, peer review can be biased toward well-known institutions, established authors, or fashionable topics (Lee et al., 2013). Reviewers may assign high scores to work that aligns with unspoken disciplinary tastes or collective values Sulik et al. (2025), while "lazily" masking these preferences in the language of standardized evaluation criteria (Purkayastha et al., 2025). As a result, much of what constitutes the tacit code of a field – its evaluative culture – remains hidden beneath the surface, shaping scientific recognition in nontransparent ways that remain hard to alter.
@@
Data: We collected data from the OpenReview API and Paper Copilot  [^2] https://papercopilot.com/, a website that aggregates peer review scores (mostly) for computer science conference submissions. Our final dataset includes metadata, peer review scores, and review comments for 26,731 papers submitted to 46 conferences including computer science, physics, medicine, and social sciences. We use peer review scores as a proxy for the paper’s perceived quality, as shown in Appendix A, due to the consistency and high confidence of judgments among reviewers.
@@
To represent each paper’s content, we follow the approach of (Zhang et al., 2025), constructing an extended abstract that integrates additional structured elements — contextual background, key ideas, methodological and theoretical details, experiments and results, and mentioned impact. We generate these extended abstracts using prompts used in (Zhang et al., 2025) with GPT-4o. This representation generation has been empirically shown to achieve high quality across evaluation metrics and to closely match human annotations. It also demonstrates strong alignment with the original text and a low rate of hallucination (Amar et al., 2023; Zhang et al., 2025), see Appendix E for fidelity experiments. We deliberately avoid feeding full texts directly to LLMs for two main reasons. First, context windows remain limited, preventing broad coverage that would enable more generalizable hypotheses. Current approaches, e.g., (Movva et al., 2025), employ reduction methods such as sparse autoencoders to compress long inputs into activated neurons for hypothesis generation, but these are typically applied only to straightforward cases (e.g., political attitudes) where activations can be reduced to obvious dimensions (e.g., gender and race). Second, prior work has shown that LLMs often struggle with long, unstructured inputs, leading to degraded performance and increased hallucination rates mid-text (Liu et al., 2024b; Yen et al., 2024). When presented with lengthy texts, attention may diffuse across irrelevant sections, impairing reasoning over core content. To address these issues, we adopt a curated and condensed input format, ensuring that the model attends to the most relevant information. Dataset and paper representations are detailed in Appendix H. Most importantly, robustness checks in the Appendix E show that LLM judgments on full papers and extended abstracts yield highly consistent results (consistency = 0.89).
@@ 3 The Framework: Prior, Posterior, and Hypothesis Generation
We adopt a hypothesis generation framework – a topic increasingly explored in the automation of scientific discovery using AI (Radensky et al., 2024; Liu et al., 2024a; Batista and Ross, 2024; Liu et al., 2025) – to update LLMs’ stereotyped priors about good science into posteriors that reflect how human evaluators and LLMs actually judge.
@@
Self-consistency Refinement: We apply the five generated hypotheses to all pairs in the dataset. For each pair and hypothesis, we ask the LLM to determine the degree to which the hypothesis explains the observed quality difference (human review scores) using the following prompt: "Analyze the content from both papers to determine whether Paper 2 is superior in the aspects identified in the hypothesis (i.e., if the comparison of content supports the hypothesis)." This type of self-confirmation explicitly reveals the inherent biases of LLMs during the judgment process. The LLM outputs two things: (1) a binary label (0 or 1) indicating whether the hypothesis explains the score difference, and (2) a confidence score (0–10) for its judgment. We use 3-fold voting and compute the final label using a confidence-weighted scheme: if the sum of confidence scores for the label "1" exceeds that of label "0", the final label is 1; otherwise, it is 0. This confidence-weighted approach has been shown to improve self-consistency and accelerate convergence to the stable answer of LLMs across a variety of QA tasks (Taubenfeld et al., 2025). Beyond improving reliability, confidence scores across multiple judgments provide a fine-grained signal that distinguishes borderline cases from clear-cut ones - an especially valuable property in our setting, where evaluating internal characteristics is inherently difficult (Bao and Teplitskiy, 2024) and confidence levels can help identify when the model is genuinely uncertain (Taubenfeld et al., 2025). We also tested majority voting, and the resulting labels are highly correlated with the confidence-weighted labels (correlation > 0.98). LLMs may exhibit position bias, meaning that when serving as judges, they tend to favor the candidate appearing in a specific position (e.g., the first) (Shi et al., 2024). We adopt standard strategies to mitigate this bias. First, we construct paper pairs with substantial quality differences; prior research has shown that position bias is the most pronounced when the gap is small (Li et al., 2024b; Shi et al., 2024). Second, we randomize the positions of Paper 1 and Paper 2 across voting while maintaining logically consistent prompts and voting aggregation. Consistency of judgments across randomized positions achieves 77% and we have more discussions in Appendix M.
@@
We are particularly interested in belief change during this process, i.e., how LLM’s prior about good science transforms into a robust posterior, after judging the actual pairwise dataset. For each generated hypothesis, when the LLM judges the pairwise dataset, we retain the proportion of cases it can explain (one pair can be explained by multiple hypotheses). We also match each hypothesis to priors (prompting: do you think they are talking about the same idea and same hypothesis?) to obtain their appearance frequency in the prior. We also conducted cross-validation experiments for the reliability of LLM labels of "match" using a heuristic dictionary-based method (locating some keywords of posterior hypotheses in priors to identify "match"), and these two methods exhibit a relatively high degree of correlation of 0.70, pp=0.001 (see Appendix K for details). The frequency of each posterior hypothesis in the prior distribution is measured by counting how often a similar hypothesis appears within each window of 20 sequential prior hypotheses across 250 simulations. We use 20-hypothesis windows to align with the posterior generation setup, where LLMs are prompted to produce 4 rounds of 5 unique hypotheses. This ensures a fair comparison between the frequency of hypotheses in the prior and posterior distributions.
@@ 4.2 Human judgments and LLMs share priors, but not posteriors
We observe notable shifts in the frequency of certain hypotheses from prior to posterior, as shown in Figure 4. Some hypotheses remain relatively stable in frequency and their explanatory power for good science (middle). In contrast, others manifest substantial change – some decrease markedly in frequency (left), while others increase significantly (right) from the prior distribution (see Table 1). This shift reflects a transition, after being exposed to the pairwise data, from emphasizing normative scientific values – such as theoretical or methodological rigor - to prioritizing more practical, communicative, and contextualization aspects, including storytelling quality and interdisciplinary relevance. When treated as a peer reviewer, the LLM initially aligns with conventional scientific ideals. In practice, however, it leans more heavily on accessible, narrative-driven criteria that appear more explanatory in distinguishing lower from higher rated papers (Figure 3 panels c and d). This suggests a divergence between explicit and implicit criteria of evaluation in science, a phenomenon widely appreciated although rarely measured in social studies of science Latour (1987).
@@
We collected human review comments and annotated them using in-context learning prompts with GPT-4o, leveraging its strong performance in social computing annotation tasks, often surpassing human experts and crowd-workers (Törnberg, 2023; Alizadeh et al., 2023; Haq et al., 2025). For each hypothesis, the annotation model was provided with three illustrative examples of reviewers not-mentioning (label 0) specific aspects in a hypothesis, along with the reviewers’ attitudes toward those aspects (positive: 1 or negative: -1, they are considered as "mentioning"). For robustness, we conducted two experiments: 1) the dictionary method produces high consistency with LLM’s "mention" label with correlation of 0.78, $p=0.000$, detailed in Appendix K. 2) Three master’s students specializing in social computing independently conducted the human annotations, achieving a high level of consistency with the LLM-generated labels (see Appendix B for details). We find human comments align with LLM’s priors in terms of the frequency of aspects they address, meaning that they articulate a similar attention distribution as LLM’s prior. In contrast, human reviewers rarely articulate their updated judgments like the LLM does. That is, if we assume the LLM mirrors human reviewing behavior, both begin from similar priors. However, while the LLM transparently updates its evaluation (posterior), human reviewers often refrain from articulating such shifts – likely due to normative expectations in academic reviewing – despite favoring unarticulated values with higher scores (we have more discussions and empirical evidence in Section 5).